OpenStudy (anonymous):

6^5x=20 round to te nearest ten- thousandth? i have most of the work down just need someone to verify if i am correct or not.

OpenStudy (anonymous):

i will show te work here that i have done.

OpenStudy (anonymous):

\[6^5x=20\] the x should be with the 5

OpenStudy (anonymous):

now the work log6^5x = log 20 (5x) log 6 = log 20 (5x) =( log 6/ log 20 ) x= (log 20/ log 6) / 5 x=.153724357 which means that x= .1537 because two doeasnt make the seven round up it stays the same.

OpenStudy (anonymous):

or did i do that wrong?

OpenStudy (phi):

ok up to (5x) log 6 = log 20 we can write this as (5 log 6) x = log20 5log6 is some number (ugly decimal) but a number. Divide both sides by this ugly number \[ \frac{(5 \log 6)x}{(5 \log 6)} = \frac{\log 20}{(5 \log 6)}\] anything divided by itself is 1 , so on the left side of the = we have \[ \frac{\cancel{(5 \log 6)}x}{\cancel{(5 \log 6)}} = \frac{\log 20}{(5 \log 6)}\] now can you finish ?

OpenStudy (anonymous):

can i put it in my calculator as 6 log 5 because i have to put the number befor the log in my calculator

OpenStudy (phi):

it is log(20) divided by (5 times log 6)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok so i got .043367667 in my calcuator rounded to the nearest ten thousandth it would be .0434 correct?

OpenStudy (phi):

try again. log20 = 1.30103 5 * log 6 = 3.890756

OpenStudy (anonymous):

ok so nw i have a large number 8.359750081 rounded to the nearest ten thousandth would be 8.3598

OpenStudy (phi):

no. I can't tell how you are getting these numbers, but they are not correct. log20 / ( 5 log6) = 1.30103/3.890756 see if you can get those numbers. the answer is about 0.3....

OpenStudy (anonymous):

i dont know how o put the equation into the calculator correctly because in order to put a log in you put the number before thelog so five log six is a terrible thing to put into the calculator for me because i dont understand how

OpenStudy (anonymous):

i only get a point three numbe rif i divide by 5 then get 0.334390003 which rounds out to .3344

OpenStudy (phi):

ok, that is the correct answer. but to do this with your calculator, do you type in 20 then log ? and get a number 1.30103

OpenStudy (anonymous):

yes

OpenStudy (phi):

then type 5 and divide and get 0.26020599

OpenStudy (anonymous):

oh ok thank you i am about to post another question which i dont have a clue how to do if ou would please help me.

OpenStudy (phi):

then type 6 log divide and this should give you the final number 0.3344 you might want to get a different calculator that is easier to use. or use google type log(20)/(5*log(6))= into the google search window